# Security Vulnerabilities in Ethereum Smart Contracts: A Systematic Analysis

### v3

@@ Abstract.
Smart Contract; Ethereum; Blockchain; Security Vulnerability
1. INTRODUCTION
Since its inception, blockchain technology has been rapidly driving the development of the digital economy with its unalterability and decentralization, but it is also facing more and more threats. Ether, as a mainstream smart contract platform, has also suffered many attacks in recent years. For example, the 2016 TheDAO incident led to a loss of \$60 million, the 2017 Parity wallet vulnerability froze 150,000 ETH, and the 2020 Coin Exchange was attacked with a loss of over \$400 million(Krichen et al., 2022). The advantage of smart contracts lies in their financial attributes, but their code vulnerabilities and tool limitations constrain their further development. Research on smart contract security is also in full swing, with Kosba et al. proposing the privacy-preserving framework Hawk, Dickerson’s team designing a parallel execution architecture to improve throughput, and the Oyente tool detecting contract vulnerabilities through symbolic execution. However, the current research work still faces many challenges, such as dynamic environment simulation and static rule detection are difficult to balance the coverage rate and false alarm rate, and formal verification tools are still insufficient in detecting the accuracy of composite attacks. Meanwhile(Chu et al., 2023), related projects at home and abroad show differentiated development paths - domestic work focuses on compliance and cross-chain interoperability, while foreign work focuses on decentralization performance.
@@ 2.1. Blockchain
Blockchain was initially intended to be the technology that drives Bitcoin, but it can do much more. It is an abbreviated expression of a set of distributed ledger technologies that can be written to record and track anything of value, including financial transactions and medical records.(Song et al., 2024) Blockchain stores information in batches, each batch is called a block, and they are connected in a line in chronological order, like a chain of blocks. The modification of block information is different from our perception. Suppose the information in a block is to be modified at this point; it does not rewrite this information, but first changes it and then stores it in a new block. Immediately afterward, the record will show that information X was changed to information Y at some point in time, a non-destructive way of tracking data that stems from the blockchain’s foundation, the time-honored financial ledger recording method. Blockchain, unlike the age-old way of keeping records, is decentralized by design; it is distributed across a huge network of computers, and this decentralization of information greatly reduces the likelihood of data tampering. Since blockchain generates data trust, this allows us the opportunity to interact with the data in real-time, and this brings us to the reason why blockchain technology is a game-changer, i.e., there is no longer a need for an intermediary. Interacting with data trusted peer-to-peer can revolutionize the way we store data, verify information, and transact with each other. Blockchain is a technology rather than a single network, and it can be implemented in a variety of ways. Some blockchains can be completely public for all to view and access, and others can be closed to a designated group of authorized users, such as companies and government agencies (Sapra et al., 2023).
2.2. Ether
@@
Ether was founded by “V-God” in 2015, as experts interested in Bitcoin could only tinker with it, not fundamentally solve its flaws, and make it too single-functional. Ether is a global, open-source, public blockchain platform, which can be simply understood as the underlying system of the Blockchain. It is like in the PC era, our underlying operating system is the Windows system(Li et al., 2021), and in the era of mobile internet, the cell phone has the Android system as well as the iOS system. These systems are the underlying technologies used by all software applications to develop their applications. Anyone can go ahead and build decentralized apps on the Ethereum platform using the blockchain running through it. The Ether system is different from Windows, Android, and Apple; it is run by nodes all over the world, so it makes Ether more secure and transparent. As an innovative platform, just as protocols are the foundation of the internet, it is not easy to predict what they are used for. The Ether project expects to do the same for finance, peer-to-peer trade, distributed governance, and human cooperation (Javaid et al., 2022). The Ethernet system architecture is shown in Figure 1.
Figure 1. The Ethernet System Architecture
2.3. Main components of Ether
@@ 2.3.1   Blocks
Block consists of three parts: Block header, Transaction list, and Uncle block, Its main use is to record the transactions generated over some time and the results of the state.
@@ 2.3.2   Accounts
Accounts use the address as their index, the public key is the basis for the address, and the last 20 bytes of the public key refer to the address. Ethernet accounts are categorized into external accounts and contract accounts(Li et al., 2020).
@@
  • Contract Account: Created by an external account. Unlike an external account, a contract account contains the contract code. As opposed to an external account, it is controlled by the contract code rather than the private key. Ether rationing, the ability to activate the code by invoking a transaction or other contract, and the fact that a contract account can only manipulate the storage it owns when it is executed are all characteristics of a contract account. The operations on the Blockchain, if they are to be carried out properly, can only be based on the transactions of the account. If a parameter is entered into the contract code, the parameter can be called on the contract code only if the contract account receives the transaction information.
@@ 2.3.3   Consensus Mechanisms
PoW.The “Witch” attack, which is very harmful to the security of the system, mainly through the control of most of the nodes in the system, the modeled in Figure 2. The way to prevent the “witch” attack is to make the creation of blocks no longer easy, so the PoW mechanism proves that the way to complete the work is not through the process but through the result (Gadekallu et al., 2022). The specific algorithm of PoW is related to the hash function. The hash function can transform any string of any length into a fixed-length hash value, characterized by being collision-free, steganographic, and with the hash value falling within a fixed interval.
Figure 2. The ”Witch Attack” Model
PoS.Since PoS can allocate time across several digital currencies owned by participants in the network, it can generally be used to prove user ownership. The general situation is similar to PoW, but the difference is that the virtual consumption of power is negligible. From the algorithmic point of view, there are two types: chain-based PoS (Gallenmüller et al., 2021)and BFT. With chain-based PoS, it gives the right to the verifier to create a new block, the most blocks to be pointed to by this new block, and this verifier is chosen by the algorithm from the set of verifiers. The further back in time, the more centralized the blocks become. For BFT, it will give some rights to the verifiers with the aim that these verifiers have the right to be able to propose blocks, and also to be able to select which block is to be the new block, they should also vote for these blocks. And in each round, a new block is elected to be added to the blockchain.
@@ 2.3.4   Transaction
Smart contracts run on Ether(Kushwaha et al., 2022a), and since Ether is Turing-complete, it can execute infinitely complex, cyclic code, to ensure that malicious code is not executed and that resources on the network are not wasted. Gas, as well as Gas Prices, were introduced. Gas is related to the computational resources consumed in a transaction. Different computational needs will have different Gas, for example, the computational resources needed to do addition, subtraction, read an article(Li et al., 2024d), or read code are different. Gas Price refers to the amount of gas consumed by one unit of Gas(Liu et al., 2024). Common Gas consumption and operations are shown in Table 1.
Table 1. Common Operations and Consumption
@@ 2.4.1   Applications of Smart Contracts
Nick Szabo first coined the term “smart contract” in 1997 (Perez and Zeadally, 2022), which simply means that he wanted to preserve contracts through a distributed ledger. Smart contracts are similar to real-life contracts, but the only difference is that they are completely digital; they are like a small program that is stored in the blockchain. As an example, Kickstarter is a large crowdfunding site that, by its very nature, is based on a third party between the project team and its backers, so both parties need to trust that Kickstarter can keep the funds(Li et al., 2025). With smart contracts, it is possible to create a similar crowdfunding system without the need for a third party, where project backers first transfer money to the smart contract, and if the project is fully funded, the smart contract automatically transfers the money to the project creator, and if no funds are available to reach the goal, the money is automatically returned to project backers. Using this technology, no one can control the money, and we trust the smart contract because it is kept on the blockchain, and thus inherits many of the blockchain’s properties, such as the fact that they are distributed and cannot be tampered with. Smart contracts can also be used in places other than just crowdfunding. Banks could use it to issue loans as well as provide automated payments, insurance companies could use it to process claims, courier companies could use it for cash on delivery, and so on. But there are only a few blockchains that support smart contracts today, and the largest of them is Ether. Additional information about the smart contract is shown in Figure 3.
@@ 2.4.2   Smart Contract Related Issues
  • In terms of Privacy. Currently, blockchain has been applied in many fields such as healthcare, finance, Internet of Things, etc., which stems from the fact that it has the characteristics of de-trust, decentralization, open autonomy, and information tampering. Unfortunately, the anonymity of the blockchain does not completely solve the privacy problem of smart contracts(Li et al., 2023c), and for this reason, Kosba et al. proposed Hawk, a framework for the development of smart contracts. In this framework, to protect the privacy of the user, the blockchain will not be displayed on the blockchain to record all the financial information. To ensure that there is no way for other users in the blockchain to view the content of the request, Zhang et al. proposed Town Crier, a trusted data entry system.
@@
  • In terms of Performance. In the existing blockchain system, smart contracts are executed sequentially. Dickerson et al. proposed a parallel execution framework based on smart contracts for multi-core and clustered architectures that are executable per second. This architecture allows different contracts to run in parallel(Li et al., 2024f), thus increasing the throughput of the system while improving the efficiency of smart contract execution.
@@
  • In terms of Security. Up to now, of the 12 security vulnerabilities that exist in Ethernet smart contracts, the most common are transaction order dependency, timestamp dependency, processing anomalies, and reentrancy dependency(Li et al., 2024a). To prevent attackers from changing transaction orders, modifying timestamps, calling reentrant functions, and triggering processing exceptions, which can affect the execution of smart contracts, Oyente was proposed as a symbolic execution tool by Luu et al(Li et al., 2017). It reviewed 19,366 smart contracts, of which 8,833 had vulnerabilities in the above areas. Chen and others invented a high-consumption burning smart contract testing tool called Gasper, which automatically finds dead code(Kong et al., 2024), computationally intensive looping operations, and so on. Using Gasper, they discovered that more than 80% of smart contracts in Ether are subject to high-consumption operations, which are the cause of denial of service.
Figure 3. Mechanisms and Applications of Smart Contracts
3. CATEGORIES
3.1. Smart Contract Vulnerability Classification and Definition
@@
In layman’s terms, a contract vulnerability is a problem that exists in the code when writing a contract, which can cause a large amount of asset loss once exploited by hackers, since money is involved. In Luu’s Making Smart Contracts Smarter article, the author describes Transaction ordering dependence, Timestamp dependence, and mishandled exceptions, and proposes an open-source detection tool(Li et al., 2024c), Oyente, which can help analyze Ethereum smart contracts and detect most of the vulnerabilities that already exist in the contracts. In A Survey of Attacks on Ethereum Smart Contracts(Zhou et al., 2022), the authors Atzei et al. systematized the security vulnerabilities in Ethereum smart contracts and classified the vulnerabilities into three categories based on the level of vulnerabilities: the Solidity layer, the EVM layer, and the Blockchain layer(Niu et al., 2024). As shown in Table 2.
Table 2. Smart Contract Vulnerability Classification
@@ 3.2.2   Principle of Integer Overflow Attack
In any programming language, it has integer types, such as int and uint. For signed integers, no matter whether int8 or int64, there will be a maximum value. int8’s upper limit value is 127, so a statement like int8 a=129 is impossible to pass in programming, but if the statement is changed to int8 a=100+29, in This time the compiler will first react to the final value of the addition, and then go to judgment, so this statement will pass in programming, and because the final result 129 is greater than 128, which leads to overflow. Here we take the 8-bit unsigned integer type as an example, the value range of uint8 is [0,255], and the form of storing in memory by bit is shown in Figure 4. Obviously, an unsigned integer, expressed in binary(Wang et al., 2023), is all 1’s and occupies exactly 8 bits, and if you add 1’s at this point, then you’ll be performing a different-or operation, which will result in all 8 bits being filled with 0’s at this point.
Figure 4. Uint8 Storage Form
3.3. Reentry Vulnerability
@@ 3.3.1   The DAO security incident
On June 18, 2016, a hacker exploited a vulnerability in The DAO to steal 3.6 million Ethereum coins from the project’s smart contract, which was worth about \$50 million at the time. This was due to a split function in that contract. The attacker used the vulnerability in this function to get this asset by using his own DAO asset to continuously separate the DAO(Wang et al., 2022) asset from the asset pool of The DAO project. So, people call this vulnerability a re-entry vulnerability.
@@ 3.3.2   Reentry Vulnerability Principle
The cause is that when the contracts make mutual calls or interactions, the external contract gets control, which will modify the data in a way that should not be there. An attacker can write a contract first, and since the address of the contract and the address of the account have the same number of bits, although there are ways to identify it, it is not distinguishable on the surface. This is shown in Figure 5. First of all, the attacker will write a contract, and their contract to fill the money, and then will go to call the withdraw function, the attacked contract with a call(Li et al., 2024b), which is an external call, at this time the attacker will get the right to call, so the attacker back to their contract, the use of rollback function (fallback) and then go to call the withdraw function, at this time back to the first step, resulting in the entry of a contract. The first step leads to a loop.
Figure 5. Reentry attack process
3.4. Denial of Service Attacks

### v4

@@ Abstract
keywords:
Smart Contract, Ethereum, Blockchain, Security Vulnerability
journal: Cyber Security and Applications
[…]
organization=Hainan University,city=Haikou, postcode=570228, state=Hainan, country=China
[…]
organization=Hainan University,city=Haikou, postcode=570228, state=Hainan, country=China
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author: Xiaoqi Li. E-mail: csxqli@ieee.org.
[…]
organization=Hainan University,city=Haikou, postcode=570228, state=Hainan, country=China
{highlights}
This paper analyzes typical security events in the Ethernet smart contract environment and reveals the security vulnerability principles behind them. We study the triggering mechanism and hazards of related vulnerabilities.
Based on the Solidity language and Remix development environment, this paper reproduces the vulnerability scenarios and proposes defense strategies.
This paper lays a theoretical foundation for the future development of smart contract protection technology through systematic categorization and case studies.
1 INTRODUCTION
Since its inception, blockchain technology has been rapidly driving the development of the digital economy with its unalterability and decentralization, but it is also facing more and more threats. Ether, as a mainstream smart contract platform, has also suffered many attacks in recent years. For example, the 2016 TheDAO incident led to a loss of \$60 million, the 2017 Parity wallet vulnerability froze 150,000 ETH, and the 2020 Coin Exchange was attacked with a loss of over \$400 million[34]. The advantage of smart contracts lies in their financial attributes, but their code vulnerabilities and tool limitations constrain their further development. Research on smart contract security is also in full swing, with Kosba et al. proposing the privacy-preserving framework Hawk, Dickerson’s team designing a parallel execution architecture to improve throughput, and the Oyente tool detecting contract vulnerabilities through symbolic execution[1]. However, the current research work still faces many challenges, such as dynamic environment simulation and static rule detection are difficult to balance the coverage rate and false alarm rate, and formal verification tools are still insufficient in detecting the accuracy of composite attacks. Meanwhile[10], related projects at home and abroad show differentiated development paths - domestic work focuses on compliance and cross-chain interoperability, while foreign work focuses on decentralization performance[3].
@@ 2.1 Blockchain
Blockchain was initially intended to be the technology that drives Bitcoin, but it can do much more. It is an abbreviated expression of a set of distributed ledger technologies that can be written to record and track anything of value, including financial transactions and medical records.[63] Blockchain stores information in batches, each batch is called a block, and they are connected in a line in chronological order, like a chain of blocks. The modification of block information is different from our perception. Suppose the information in a block is to be modified at this point; it does not rewrite this information, but first changes it and then stores it in a new block. Immediately afterward, the record will show that information X was changed to information Y at some point in time, a non-destructive way of tracking data that stems from the blockchain’s foundation, the time-honored financial ledger recording method. Blockchain, unlike the age-old way of keeping records[9], is decentralized by design; it is distributed across a huge network of computers, and this decentralization of information greatly reduces the likelihood of data tampering[4]. Since blockchain generates data trust, this allows us the opportunity to interact with the data in real-time, and this brings us to the reason why blockchain technology is a game-changer, i.e., there is no longer a need for an intermediary. Interacting with data trusted peer-to-peer can revolutionize the way we store data, verify information, and transact with each other. Blockchain is a technology rather than a single network, and it can be implemented in a variety of ways. Some blockchains can be completely public for all to view and access, and others can be closed to a designated group of authorized users, such as companies and government agencies [59].
2.2 Ether
@@
Ether was founded by “V-God” in 2015, as experts interested in Bitcoin could only tinker with it, not fundamentally solve its flaws, and make it too single-functional[8]. Ether is a global, open-source, public blockchain platform, which can be simply understood as the underlying system of the Blockchain. It is like in the PC era, our underlying operating system is the Windows system[43], and in the era of mobile internet[5], the cell phone has the Android system as well as the iOS system. These systems are the underlying technologies used by all software applications to develop their applications. Anyone can go ahead and build decentralized apps on the Ethereum platform using the blockchain running through it. The Ether system is different from Windows, Android, and Apple; it is run by nodes all over the world, so it makes Ether more secure and transparent. As an innovative platform, just as protocols are the foundation of the internet[7], it is not easy to predict what they are used for. The Ether project expects to do the same for finance, peer-to-peer trade, distributed governance, and human cooperation [29]. The Ethernet system architecture is shown in Figure 1.
Figure 1: The Ethernet System Architecture
2.3 Main components of Ether
@@ 2.3.1 Blocks
Block consists of three parts: Block header, Transaction list, and Uncle block, Its main use is to record the transactions generated over some time and the results of the state[6].
  • 1.
@@ 2.3.2 Accounts
Accounts use the address as their index, the public key is the basis for the address, and the last 20 bytes of the public key refer to the address. Ethernet accounts are categorized into external accounts and contract accounts[13].
  • 1.
@@
  • 2.
    Contract Account: Created by an external account. Unlike an external account, a contract account contains the contract code. As opposed to an external account, it is controlled by the contract code rather than the private key. Ether rationing, the ability to activate the code by invoking a transaction or other contract, and the fact that a contract account can only manipulate the storage it owns when it is executed are all characteristics of a contract account. The operations on the Blockchain, if they are to be carried out properly, can only be based on the transactions of the account. If a parameter is entered into the contract code, the parameter can be called on the contract code only if the contract account receives the transaction information[14].
@@ 2.3.3 Consensus Mechanisms
PoW.The “Witch” attack, which is very harmful to the security of the system, mainly through the control of most of the nodes in the system, the modeled in Figure 2. The way to prevent the “witch” attack is to make the creation of blocks no longer easy, so the PoW mechanism proves that the way to complete the work is not through the process but through the result [20]. The specific algorithm of PoW is related to the hash function. The hash function can transform any string of any length into a fixed-length hash value, characterized by being collision-free, steganographic, and with the hash value falling within a fixed interval.
Figure 2: The ”Witch Attack” Model
PoS.Since PoS can allocate time across several digital currencies owned by participants in the network, it can generally be used to prove user ownership. The general situation is similar to PoW, but the difference is that the virtual consumption of power is negligible. From the algorithmic point of view, there are two types[15]: chain-based PoS [21]and BFT. With chain-based PoS, it gives the right to the verifier to create a new block, the most blocks to be pointed to by this new block, and this verifier is chosen by the algorithm from the set of verifiers. The further back in time, the more centralized the blocks become. For BFT, it will give some rights to the verifiers with the aim that these verifiers have the right to be able to propose blocks, and also to be able to select which block is to be the new block, they should also vote for these blocks. And in each round, a new block is elected to be added to the blockchain.
@@ 2.3.4 Transaction
Smart contracts run on Ether[35], and since Ether is Turing-complete, it can execute infinitely complex, cyclic code, to ensure that malicious code is not executed and that resources on the network are not wasted. Gas, as well as Gas Prices, were introduced. Gas is related to the computational resources consumed in a transaction. Different computational needs will have different Gas, for example, the computational resources needed to do addition, subtraction, read an article, or read code are different. Gas Price refers to the amount of gas consumed by one unit of Gas. Common Gas consumption and operations are shown in Table 1.
Table 1: Common Operations and Consumption
@@ 2.4.1 Applications of Smart Contracts
Nick Szabo first coined the term “smart contract” in 1997 [53], which simply means that he wanted to preserve contracts through a distributed ledger. Smart contracts are similar to real-life contracts, but the only difference is that they are completely digital; they are like a small program that is stored in the blockchain[16]. As an example, Kickstarter is a large crowdfunding site that, by its very nature, is based on a third party between the project team and its backers, so both parties need to trust that Kickstarter can keep the funds. With smart contracts, it is possible to create a similar crowdfunding system without the need for a third party, where project backers first transfer money to the smart contract, and if the project is fully funded, the smart contract automatically transfers the money to the project creator, and if no funds are available to reach the goal, the money is automatically returned to project backers. Using this technology, no one can control the money, and we trust the smart contract because it is kept on the blockchain, and thus inherits many of the blockchain’s properties, such as the fact that they are distributed and cannot be tampered with. Smart contracts can also be used in places other than just crowdfunding. Banks could use it to issue loans as well as provide automated payments, insurance companies could use it to process claims, courier companies could use it for cash on delivery[19], and so on. But there are only a few blockchains that support smart contracts today, and the largest of them is Ether. Additional information about the smart contract is shown in Figure 3.
@@ 2.4.2 Smart Contract Related Issues
    In terms of Privacy. Currently, blockchain has been applied in many fields such as healthcare, finance, Internet of Things, etc., which stems from the fact that it has the characteristics of de-trust, decentralization, open autonomy, and information tampering. Unfortunately, the anonymity of the blockchain does not completely solve the privacy problem of smart contracts, and for this reason[22], Kosba et al. proposed Hawk, a framework for the development of smart contracts. In this framework, to protect the privacy of the user, the blockchain will not be displayed on the blockchain to record all the financial information. To ensure that there is no way for other users in the blockchain to view the content of the request, Zhang et al. proposed Town Crier, a trusted data entry system.
  • 2.
@@
    In terms of Performance. In the existing blockchain system, smart contracts are executed sequentially. Dickerson et al. proposed a parallel execution framework based on smart contracts for multi-core and clustered architectures that are executable per second. This architecture allows different contracts to run in parallel[24], thus increasing the throughput of the system while improving the efficiency of smart contract execution.
  • 3.
@@
    In terms of Security. Up to now, of the 12 security vulnerabilities that exist in Ethernet smart contracts, the most common are transaction order dependency, timestamp dependency, processing anomalies, and reentrancy dependency. To prevent attackers from changing transaction orders[25], modifying timestamps, calling reentrant functions, and triggering processing exceptions, which can affect the execution of smart contracts, Oyente was proposed as a symbolic execution tool by Luu et al. It reviewed 19,366 smart contracts, of which 8,833 had vulnerabilities in the above areas. Chen and others invented a high-consumption burning smart contract testing tool called Gasper, which automatically finds dead code[32], computationally intensive looping operations, and so on. Using Gasper, they discovered that more than 80% of smart contracts in Ether are subject to high-consumption operations, which are the cause of denial of service.
Figure 3: Mechanisms and Applications of Smart Contracts
3 CATEGORIES
3.1 Smart Contract Vulnerability Classification and Definition
@@
In layman’s terms, a contract vulnerability is a problem that exists in the code when writing a contract[26], which can cause a large amount of asset loss once exploited by hackers, since money is involved. In Luu’s Making Smart Contracts Smarter article, the author describes Transaction ordering dependence, Timestamp dependence, and mishandled exceptions, and proposes an open-source detection tool[27], Oyente, which can help analyze Ethereum smart contracts and detect most of the vulnerabilities that already exist in the contracts[30]. In A Survey of Attacks on Ethereum Smart Contracts[73], the authors Atzei et al. systematized the security vulnerabilities in Ethereum smart contracts and classified the vulnerabilities into three categories based on the level of vulnerabilities: the Solidity layer, the EVM layer[31], and the Blockchain layer. As shown in Table 2.
Table 2: Smart Contract Vulnerability Classification
@@ 3.2.2 Principle of Integer Overflow Attack
In any programming language, it has integer types, such as int and uint. For signed integers, no matter whether int8 or int64, there will be a maximum value. int8’s upper limit value is 127, so a statement like int8 a=129 is impossible to pass in programming, but if the statement is changed to int8 a=100+29, in This time the compiler will first react to the final value of the addition, and then go to judgment, so this statement will pass in programming, and because the final result 129 is greater than 128, which leads to overflow. Here we take the 8-bit unsigned integer type as an example, the value range of uint8 is [0,255], and the form of storing in memory by bit is shown in Figure 4. Obviously, an unsigned integer, expressed in binary[69], is all 1’s and occupies exactly 8 bits, and if you add 1’s at this point, then you’ll be performing a different-or operation, which will result in all 8 bits being filled with 0’s at this point.
Figure 4: Uint8 Storage Form
3.3 Reentry Vulnerability
@@ 3.3.1 The DAO security incident
On June 18, 2016, a hacker exploited a vulnerability in The DAO to steal 3.6 million Ethereum coins from the project’s smart contract, which was worth about \$50 million at the time. This was due to a split function in that contract. The attacker used the vulnerability in this function to get this asset by using his own DAO asset to continuously separate the DAO[70] asset from the asset pool of The DAO project. So, people call this vulnerability a re-entry vulnerability.
@@ 3.3.2 Reentry Vulnerability Principle
The cause is that when the contracts make mutual calls or interactions[41], the external contract gets control, which will modify the data in a way that should not be there[44]. An attacker can write a contract first, and since the address of the contract and the address of the account have the same number of bits, although there are ways to identify it, it is not distinguishable on the surface. This is shown in Figure 5. First of all, the attacker will write a contract, and their contract to fill the money, and then will go to call the withdraw function, the attacked contract with a call, which is an external call, at this time the attacker will get the right to call, so the attacker back to their contract, the use of rollback function (fallback) and then go to call the withdraw function, at this time back to the first step, resulting in the entry of a contract. The first step leads to a loop.
Figure 5: Reentry attack process
3.4 Denial of Service Attacks